OpenStudy (anonymous):

Convert the polar equation to rectangular form r= 3sec theta/5sec theta+5

OpenStudy (surry99):

Do you know the transformations for x, y and r in terms of polar coordinates?

OpenStudy (anonymous):

Yes, x=rcos(theta), y=rsin(theta) and r=sqareroot of x^2+y^2

OpenStudy (surry99):

Excellent so: r = (3/cos(theta))/(5/cos(theta) + 5)

OpenStudy (surry99):

I left the r for a reason....

OpenStudy (anonymous):

Okay, that makes sense so far.

OpenStudy (surry99):

now you can sub for cos(theta)....in terms of x and r

OpenStudy (anonymous):

How exactly would you do that?

OpenStudy (surry99):

x=rcos(theta) therfore cos(theta) = x/r

OpenStudy (surry99):

therefore...

OpenStudy (anonymous):

I'm not sure how to move from that step...

OpenStudy (surry99):

Try to rewrite this now r = (3/cos(theta))/(5/cos(theta) + 5) with cos(theta) = x/r ( leave the r on the left for now)

OpenStudy (anonymous):

So r= (3/(x/r))/(5/(x/r) + 5)?

OpenStudy (surry99):

yes..I would simplify the right hand side as much as possible then you can sub for r... r = (x^2 + y^2)^1/2

OpenStudy (surry99):

Once you do that...it should all be just a function of x and y.

OpenStudy (anonymous):

Could you also simplify in the beginning and start off with r=5cos(theta)+5/3cos(theta)?

OpenStudy (surry99):

hang on...

OpenStudy (surry99):

When I simply the original expression r = (3/cos(theta))/(5/cos(theta) + 5) I get r = 3/(5 + 5cos(theta)

OpenStudy (anonymous):

Oh okay. I still don't understand how you would substitute for r in the initial equation

OpenStudy (surry99):

r = (x^2 + y^2)^1/2

OpenStudy (anonymous):

Would that be the square root ?

OpenStudy (surry99):

yes

OpenStudy (anonymous):

What's the simplest the right side can get?

OpenStudy (anonymous):

Can you simplify it down to 3cos(theta)+15/5cos(theta)

OpenStudy (surry99):

hang on

OpenStudy (surry99):

I am going to leave it to you to work it out. If you want to post your final answer I would be happy to look at it.

OpenStudy (anonymous):

One quick thing, how do I get rid of the square roots?

OpenStudy (surry99):

Square both sides.